# Multiple merger coalescent inference of effective population size

### v1

@@ 1 Introduction
The $\Lambda-$coalescent is a multiple merger coalescent model (MMC) proposed as a generalization to binary coalescent processes (Pitman, 1999; Sagitov, 1999). The $\Lambda-$coalescent includes Kingman’s coalescent, the Beta-coalescent (Berestycki et al., 2007) and the Psi-coalescent (Eldon and Wakeley, 2006) as special cases. In particular, the Beta-coalescent is a one-parameter model that was shown to be the ancestral limit of a sample obtained from the Cannings population model with a heavy tail offspring distribution and constant population size (Schweinsberg, 2003; Berestycki, 2009). In Hoscheit and Pybus (2019), the authors extended the Beta-coalescent model to allow variable effective population sizes and heterochronous sampling, i.e. when tips have different dates. The effective population size under Kingman’s coalescent can be interpreted as the population size under the Wright-Fisher model that has the same genetic drift as the population under study (Wakeley, 2009). Under the Beta-coalescent, the same interpretation holds for a population that evolves under the Cannings population model with a specific offspring distribution. See Eldon (2020) for a recent review of MMC models and (Korfmann et al., 2024) for modeling recombining genealogies with MMC.
@@
Most applications of the $\Lambda-$coalescent aimed to distinguish between population growth, such as exponential growth, and multiple merger genealogy from summaries of molecular data including the site frequency spectrum (SFS) (Eldon et al., 2015; Koskela, 2018; Matuszewski et al., 2018; Koskela and Berenguer, 2019; Freund and Siri-Jégousse, 2021). A recently proposed method estimates the base measure $\Lambda$ that describes the coalescent rates from the SFS (Pina et al., 2023). However all these methods assume constant effective population sizes and the infinite-alleles mutation model. Instead, a method for inferring effective population sizes under the Beta-coalescent from a given multiple merger genealogy is proposed in Hoscheit and Pybus (2019). Although this method ignores genealogical uncertainty, it opens the door to applications in which an estimated genealogy is obtained under any mutation model (for example via maximum likelihood estimation). The authors propose an extension of the classic skyline plot for inferring the effective population size (Ho and Shapiro, 2011), however there is no joint inference of effective population size and the characteristic measure that describes coalescent rates.
@@
In this manuscript, we propose a method for joint inference of the effective population size $N_{e}(t)$ and the characteristic coalescent measure under the Beta-coalescent. We expand upon the work of (Palacios and Minin, 2012; Lan et al., 2015), and develop a Bayesian nonparametric phylodynamic approach that relies on Gaussian Markov random field priors on $N_{e}(t)$. We evaluate the performance of proposed methods on simulations and apply them to analyze two infectious diseases. Finally, we re-analyze the reproductive skew hypothesis in Japanese sardine populations using our methods (Niwa et al., 2016).
@@ 2 Background on $\Lambda-$coalescent
The standard Kingman’s $n$-coalescent is a backward-in-time Markov jump chain on binary partitions of $[n]=\{1,\ldots,n\}$ whose full realization is a genealogy of $n$ individuals. The process starts with $n$ singleton lineages at time 0. At each step, two lineages are chosen uniformly at random to coalesce, continuing until there is a single lineage at the root (Kingman, 1982). The coalescent holding times are exponentially distributed with rate $\binom{A(t)}{2}$, when there are $A(t)$ ancestral lineages at time $t$. Kingman’s coalescent has also been extended to incorporate variable population size, also termed the effective population size $N_{e}(t)$ (Slatkin, 2001), and samples collected at different times (heterochronous sampling) (Felsenstein and Rodrigo, 1999). Here, $N_{e}(t)$ is interpreted as the size of an ideal population that exhibits the same level of genetic drift under a standard Wright-Fisher population model (Wakeley and Sargsyan, 2009), and can be thought of as a relative measure of genetic diversity over time.
The $\Lambda-$coalescent is a generalization of the standard coalescent that allows for multiple mergers (Sagitov, 1999; Pitman, 1999). The process is a Markov jump chain on partitions of $[n]=\{1,\ldots,n\}$ whose full realization is a multiple merger genealogy of $n$ individuals. The process starts with $n$ singleton lineages at time 0. At each step, two or more lineages are chosen to coalesce, continuing until there is a single lineage at the root (Figure 2). Here, $k\geq 2$ lineages merge at rate $\lambda_{A(t),k}$, where $A(t)$ is the number of extant lineages at time $t$. The coalescent holding times are exponentially distributed with rate $\lambda_{A(t)}=\sum_{k=2}^{A(t)}\binom{A(t)}{k}\lambda_{A(t),k}$. Conditional on a coalescent event at time $t$ when there are $A(t)$ lineages, the distribution of the block (or multiple merger) size is
$\mathbb{P}(K=k)=\frac{\binom{A(t)}{k}\lambda_{A(t),k}}{\sum_{i=2}^{A(t)}\;\;\binom{A(t)}{i}\lambda_{A(t),i}},2\leq k\leq A(t).$
@@
where $\Lambda$ is a measure on $[0,1]$. The total coalescent rate when there are $b$ lineages is defined as $\lambda_{b}=\sum_{k=2}^{b}\binom{b}{k}\lambda_{b,k}$.
Pitman (1999) showed that any simple Markovian MMC process that satisfies: exchangeability of the lineages and consistency of the merger rates defined by $\lambda_{b,k}=\lambda_{b+1,k}+\lambda_{b+1,k+1}$ must be the $\Lambda-$coalescent. The intuition behind the consistency condition is as follows. Suppose we have a specific $k$-block of lineages out of $b$ lineages. If there are $b+1$ total lineages, then the extra lineage can coalesce with the $k$-block at rate $\lambda_{b+1,k+1}$ or the extra lineage does not coalesce with the $k$-block at rate $\lambda_{b+1,k}$, and so the additive property must hold. The $\Lambda-$coalescent can also be constructed as a Poisson point process on $(0,1]\times(0,\infty)$ with intensity measure $\Lambda(dx)x^{-2}\otimes dt$; see Berestycki (2009) for a detailed theoretical construction. We can interpret points $(x_{i},t_{i})$ from this Poisson point process as flipping a coin with probability $x_{i}$ of heads for each block and merging all blocks that are heads at time $t_{i}$.
@@
We will specifically focus on the Beta$(2-\alpha,\alpha)-$coalescent with $\alpha\in(1,2]$ because it has many desirable properties. First, it converges to Kingman’s-coalescent as $\alpha\to 2$ from below. The process comes down from infinity for $\alpha\in(1,2]$, meaning that the time to the most recent common ancestor is almost surely finite. For this reason, we will only consider $\alpha\in(1,2]$. It is also the limiting distribution of a Cannings population model with a heavy tail offspring distribution and relevant for studying populations with sweepstakes reproduction. It is also relevant for studying infectious diseases. Hoscheit and Pybus (2019) showed that the $\alpha$ parameter in the Beta-coalescent model is indicative of the level of superspreading in an infectious population.
The definition of the base measure $\Lambda$, and in particular the value of $\alpha$ in the Beta-coalescent, affects both the shape of the tree topology and the distribution of the coalescent times. To see this, in Figure 1(a) we plot the average block size for different values of $\alpha$ and number of tips $n$. The average is calculated from 500 simulated trees for each value of $\alpha$. We see the average block size decreases to 2, the binary tree, as $\alpha\to 2$. The average block size also decreases with $n$, since trees with more tips are expected to have larger block sizes. In Figure 1(b) we show the total coalescent rate for different values of $\alpha$ and lineages $b$, divided by $\binom{b}{2}$, the total rate under Kingman’s coalescent. The coalescent rate increases exponentially with $\alpha$. The smaller the value of $\alpha$, the smaller the coalescent rate resulting in genealogies with longer branch lengths than with Kingman’s coalescent. This effect is similar to the one obtained from a Kingman’s coalescent model with larger effective population size (Section 2.1, Eq 3)
@@
Figure 1: The effect of $\alpha$ in tree topology and coalescent times. (a): Average block size for different values of $\alpha$ and number of tips. (b): Scaled coalescent rates when there are $b$ lineages for different values of $\alpha$, scaled by $\binom{b}{2}$, the rate under Kingman’s coalescent.
@@ 2.1 Heterochronous $\Lambda$-coalescent with variable $N_{e}(t)$
The $\Lambda-$coalescent was recently extended to accommodate heterochronous sampling and variable effective population size (Hoscheit and Pybus, 2019). Accounting for heterochronous sampling, that is considering the samples collection dates, is important for rapidly evolving organisms as it restricts the coalescent time distribution. Here, we assume that a rooted and timed multifurcating genealogy of $n$ haploid samples is available to us, for example, the estimated genealogy obtained via maximum likelihood from a set of $n$ observed molecular sequences. We further assume that sequences, at the tips of the genealogy, are collected at $L$ different sampling times and that the sampling process is independent of the underlying population process. Let $\boldsymbol{n}=(n_{\ell})_{\ell=1:L}$ denote the number of samples collected at respective sampling times $\boldsymbol{s}=(s_{\ell})_{\ell=1:L}$, with $s_{1}=0$, $s_{j-1}<s_{j}$ for $j=2,\dots,L$, and $n=\sum_{j=1}^{L}n_{j}$ is the total number of samples. In the genealogy, tuples of lineages merge backward in time into a common ancestor at coalescent times denoted by $\boldsymbol{t}=(t_{1},\ldots,t_{K})$ (see Figure 2).
@@
Let $\boldsymbol{m}=(m_{1},\ldots,m_{k})$ denote the corresponding vector of block sizes at each coalescent time. The rate at which pairs of lineages coalesce depends on the number of lineages, the $\Lambda$ measure, and the effective population size $(N_{e}(t))_{t\geq 0}:=N_{e}(t)$. Finally, let the number of extant lineages at time $t$ be given by $A(t)=\sum^{L}_{i=1}n_{i}\mathds{1}(s_{i}<t)-\sum^{K}_{k=1}\mathds{1}(t_{k}<t)$. Under this model, the density of observed statistics $\boldsymbol{m}$ and $\boldsymbol{t}$ is:
@@
In this manuscript, we are interested in estimating the effective population size $N_{e}(t)$ and the $\alpha$ parameter in the Beta-coalescent. Proposition 1 in Appendix 7.1 states that when there are $A(t)=b$ lineages, the total Lambda coalescent factor can be approximated as follows:
@@
$\log\left(\int^{t}_{0}\frac{\lambda_{b}du}{N_{e}(u)}\right)\approx\log(b-1)+(\alpha-1)\log(b/2)+\log\left(\int^{t}_{0}\frac{du}{N_{e}(u)}\right).$
@@ 2.2 Estimation of $N_{e}(t)$ using a GMRF prior
and $(\gamma_{1},\ldots,\gamma_{D-1})\sim MVN(0,(\tau Q)^{-1})$, $\tau$ is the precision parameter with Gamma prior $Gamma(0.001,0.001)$ and $Q$ is the corresponding inverse covariance kernel of a random walk with boundary correction. The integral in the exponent of Eq. 2 is then approximated by its Riemann sum. We adapted the implementation for binary trees in phylodyn (Karcher et al., 2017) to the MMC with likelihood Eq. 2. Given $\mathbf{m,s,t,n}$ (or equivalently the genealogy), and $\alpha$, our implemented R function BNPR_Lambda() estimates posterior mean and 95% BCI of $N_{e}(t)$ by an Integrated Nested Laplace Approximation (INLA) (Rue et al., 2009).

### v2

@@ 1 Introduction
The $\Lambda$-coalescent is a multiple merger coalescent model (MMC) proposed as a generalization to binary coalescent processes (Pitman, 1999; Sagitov, 1999). The $\Lambda$-coalescent includes Kingman’s coalescent, the Beta-coalescent (Berestycki et al., 2007) and the Psi-coalescent (Eldon and Wakeley, 2006) as special cases. In particular, the Beta-coalescent is a one-parameter model that was shown to be the ancestral limit of a sample obtained from the Cannings population model with a heavy tail offspring distribution and constant population size (Schweinsberg, 2003; Berestycki, 2009). Hoscheit and Pybus (2019) extended the Beta-coalescent model to allow variable effective population sizes and heterochronous sampling, i.e. when tips have different dates. The effective population size under Kingman’s coalescent can be interpreted as the population size under the Wright-Fisher model that has the same genetic drift as the population under study (Wakeley, 2009). Under the Beta-coalescent, a similar interpretation holds for a population that evolves under the Cannings population model with a specific offspring distribution. See Eldon (2020) for a recent review of MMC models and Korfmann et al. (2024) for modeling recombining genealogies with MMC.
@@
Most applications of the $\Lambda$-coalescent aimed to distinguish between population growth, such as exponential growth, and multiple merger genealogy from summaries of molecular data including the site frequency spectrum (SFS) (Eldon et al., 2015; Koskela, 2018; Matuszewski et al., 2018; Koskela and Berenguer, 2019; Freund and Siri-Jégousse, 2021). A recently proposed method estimates the base measure $\Lambda$ that describes the coalescent rates from the SFS (Pina et al., 2023). However all these methods assume constant effective population size or a fixed growth model and the infinite-alleles mutation model. Instead, a method for inferring effective population sizes under the Beta-coalescent from a given multiple merger genealogy is proposed in Hoscheit and Pybus (2019) . Although this method ignores genealogical uncertainty, it opens the door to applications in which an estimated genealogy is obtained under any mutation model (for example via maximum likelihood estimation). The authors propose an extension of the classic skyline plot (Ho and Shapiro, 2011) for inferring the effective population size, however there is no joint inference of effective population size and the characteristic measure that describes coalescent rates.
@@
In this manuscript, we propose a method for joint inference of the effective population size $N_{e}(t)$ and the characteristic coalescent measure under the Beta-coalescent. We expand upon the work of Palacios and Minin (2012); Lan et al. (2015), and develop a Bayesian nonparametric phylodynamic approach that relies on Gaussian Markov random field priors on $N_{e}(t)$. We evaluate the performance of proposed methods on simulations and apply them to analyze two infectious diseases. Finally, we re-analyze the reproductive skew hypothesis in Japanese sardine populations (Niwa et al., 2016) using our methods. Those who are interested in examining case studies can focus on Sections 4 and 5.
@@ 2 Background on $\Lambda$-coalescent
The standard Kingman’s $n$-coalescent is a backward-in-time Markov jump chain on binary partitions of $[n]=\{1,\ldots,n\}$ whose full realization is a genealogy of $n$ individuals. The process starts with $n$ singleton lineages at time 0. At each step, two lineages are chosen uniformly at random to coalesce, continuing until there is a single lineage at the root (Kingman, 1982). The coalescent holding times are exponentially distributed with rate $\binom{A(t)}{2}$, when there are $A(t)$ ancestral lineages at time $t$. Kingman’s coalescent has also been extended to incorporate variable population size, also termed the effective population size $N_{e}(t)$ (Slatkin, 2001), and samples collected at different times (heterochronous sampling) (Felsenstein and Rodrigo, 1999). Here, $N_{e}(t)$ is interpreted as the size of an ideal population that exhibits the same level of genetic drift under a standard Wright-Fisher population model (Wakeley and Sargsyan, 2009), and can be thought of as a relative measure of genetic diversity over time. The effective population size affects the coalescent times: larger $N_{e}(t)$ implies smaller coalescent rates and longer time until coalescence.
The $\Lambda$-coalescent is a generalization of the standard coalescent that allows for multiple mergers (Sagitov, 1999; Pitman, 1999). The process is a Markov jump chain on partitions of $[n]=\{1,\ldots,n\}$ whose full realization is a multiple merger genealogy of $n$ individuals. The process starts with $n$ singleton lineages at time 0. At each step, two or more lineages are chosen to coalesce, continuing until there is a single lineage at the root (Figure 2). Here, $k\geq 2$ lineages merge at rate $\lambda_{A(t),k}$, where $A(t)$ is the number of extant lineages at time $t$. The coalescent holding times are exponentially distributed with rate $\lambda_{A(t)}=\sum_{k=2}^{A(t)}\binom{A(t)}{k}\lambda_{A(t),k}$, also called the total coalescent rate. Conditional on a coalescent event at time $t$ when there are $A(t)$ lineages, the distribution of the block (or multiple merger) size is
$\mathbb{P}(X=k)=\frac{\binom{A(t)}{k}\lambda_{A(t),k}}{\sum_{i=2}^{A(t)}\;\;\binom{A(t)}{i}\lambda_{A(t),i}},2\leq k\leq A(t).$ (1)
@@
where $\Lambda$ is a measure on $[0,1]$.
Pitman (1999) show that any simple Markovian MMC process that satisfies: exchangeability of the lineages and consistency of the merger rates defined by $\lambda_{b,k}=\lambda_{b+1,k}+\lambda_{b+1,k+1}$ must be the $\Lambda$-coalescent. The intuition behind the consistency condition is as follows. Suppose we have a specific $k$-block of lineages out of $b$ lineages. If there are $b+1$ total lineages, then the extra lineage can coalesce with the $k$-block at rate $\lambda_{b+1,k+1}$ or the extra lineage does not coalesce with the $k$-block at rate $\lambda_{b+1,k}$, and so the additive property must hold. The $\Lambda$-coalescent can also be constructed as a Poisson point process on $(0,1]\times(0,\infty)$ with intensity measure $\Lambda(dx)x^{-2}\otimes dt$; see Berestycki (2009) for a detailed theoretical construction. We can interpret points $(x_{i},t_{i})$ from this Poisson point process as flipping a coin with probability $x_{i}$ of heads for each block and merging all blocks that are heads at time $t_{i}$.
@@
The definition of the base measure $\Lambda$, and in particular the value of $\alpha$ in the Beta-coalescent, affects both the shape of the tree topology and the distribution of the coalescent times. To see this, in Figure 1(a) we plot the average block size for different values of $\alpha$ and number of tips $n$. We simulate 500 trees for each value of $\alpha$ and $n$, and then compute the average block size across all trees and multiple merger events. We see the average block size decreases to 2, the binary tree, as $\alpha\to 2$. The average block size also decreases with $n$, since trees with more tips are expected to have larger block sizes. In Figure 1(b) we show the total coalescent rate for different values of $\alpha$ and lineages $b$, divided by $\binom{b}{2}$, the total rate under Kingman’s coalescent. The coalescent rate increases exponentially with $\alpha$. The smaller the value of $\alpha$, the smaller the coalescent rate resulting in genealogies with longer branch lengths than with Kingman’s coalescent.
@@
Figure 1: The effect of $\alpha$ in tree topology and coalescent times. (a): Average block size for different values of $\alpha$ and number of tips. (b): Total coalescent rates $\lambda_{b}=\sum_{k=2}^{b}\binom{b}{k}\lambda_{b,k}$ when there are $b$ lineages for different values of $\alpha$, scaled by $\binom{b}{2}$, the rate under Kingman’s coalescent.
@@ 2.1 Heterochronous $\Lambda$-coalescent with variable $N_{e}(t)$
The $\Lambda$-coalescent was recently extended to accommodate heterochronous sampling and variable effective population size (Hoscheit and Pybus, 2019). It is important to account for heterochronous sampling, especially for rapidly evolving organisms: the coalescent time distribution will be restricted since samples cannot coalesce before they have been sampled. Here, we assume that a rooted and timed multifurcating genealogy of $n$ haploid samples is available to us, for example, the estimated genealogy obtained via maximum likelihood from a set of $n$ observed molecular sequences. We further assume that sequences, at the tips of the genealogy, are collected at $L$ different sampling times and that the sampling process is independent of the underlying population process. Let $\boldsymbol{n}=(n_{\ell})_{\ell=1:L}$ denote the number of samples collected at respective sampling times $\boldsymbol{s}=(s_{\ell})_{\ell=1:L}$, with $s_{1}=0$, $s_{j-1}<s_{j}$ for $j=2,\dots,L$, and $n=\sum_{j=1}^{L}n_{j}$ is the total number of samples. In the genealogy, tuples of lineages merge backward in time into a common ancestor at coalescent times denoted by $\boldsymbol{t}=(t_{1},\ldots,t_{K})$, where $K$ is the step at which the most recent common ancestor (MRCA) is reached (see Figure 2).
@@
Let $\boldsymbol{m}=(m_{1},\ldots,m_{k})$ denote the corresponding vector of block sizes at each coalescent time. The rate at which pairs of lineages coalesce depends on the number of lineages, the $\Lambda$ measure, and the effective population size $(N_{e}(t))_{t\geq 0}:=N_{e}(t)$. Finally, let the number of extant lineages at time $t$ be given by $A(t)=\sum^{L}_{i=1}n_{i}\mathds{1}(s_{i}<t)-\sum^{K}_{k=1}(m_{k}-1)\mathds{1}(t_{k}<t)$. Under this model, the density of observed statistics $\boldsymbol{m}$ and $\boldsymbol{t}$ is:
@@
The density corresponds to the density of a point process of coalescent events with rate $\lambda_{A(t_{k})}=\sum_{j=2}^{A(t_{k})}\binom{A(t_{k})}{j}\lambda_{A(t_{k}),j}$ when there are $A(t_{k})$ lineages at time $t_{k}$, and the probability of $K$ coalescent events each being of size $m_{k}$:
$\prod_{k=1}^{K}\frac{\binom{A(t_{k})}{m_{k}}\lambda_{A(t_{k}),m_{k}}}{\lambda_{A(t_{k})}}.$ (4)
@@
In this manuscript, we are interested in estimating the effective population size $N_{e}(t)$ and the $\alpha$ parameter in the Beta$(2-\alpha,\alpha)-$coalescent with $\alpha\in(0,2]$ because it has desirable properties. As mentioned above, it is the limiting distribution of a specific Cannings population model. This model also converges to Kingman’s coalescent model as $\alpha\to 2$ from below. Hoscheit and Pybus (2019) showed that the $\alpha$ parameter in the Beta-coalescent model is indicative of the level of superspreading in an infectious population. The mathematical basis of the Beta-coalescent for $\alpha\in(0,1]$ does not change, so for the sake of implementation, we allow $\alpha\in(0,2]$ in parameter estimation. However, most simulation results will focus on Beta-coalescent trees generated with $\alpha>1$.
Proposition 1 in Appendix 7.1 states that when there are $A(t)=b$ lineages, the total Lambda coalescent factor can be approximated as follows:
@@
$\log\left(\int^{t}_{0}\frac{\lambda_{b}du}{N_{e}(u)}\right)\approx(\alpha-1)\log(b/2)+\log(b-1)+\log\left(\int^{t}_{0}\frac{du}{N_{e}(u)}\right).$
@@
Interpretation of $N_{e}(t)$ requires knowledge of the population’s variance in the number of offspring. In particular, under the Cannings population model, with variance of the number of offspring $c_{N}$, and assuming $c_{N}$ converges to 0 as $N\to\infty$, and the offspring distribution follows $\mathbb{P}(\nu_{1}>x)\sim Cx^{-\alpha}$ for $21\leq\alpha<$ 2 (Schweinsberg, 2003; Freund, 2020), then the coalescent process converges to the Beta coalescent with
$N_{e}(t)=\lim_{N\to\infty}\frac{N(\lceil x/c_{N}\rceil)}{1/c_{N}}.$
In general, we assume populations whose effective population size is scaled in units of $1/c_{N}$ generations. We note that this factor depends on $\alpha$ in the case of the Cannings model, however, as long as $c_{N}$ remains constant over time, we can jointly estimate $N_{e}(t)$ and $\alpha$ for the Beta-coalescent and solve for $N(t)$ if $c_{N}$ is known. Alternatively, the exact parametric form of $c_{N}$ could be incorporated directly in Eq. 3.
@@ 2.2 Estimation of $N_{e}(t)$ using a GMRF prior
and $(\gamma_{1},\ldots,\gamma_{D-1})\sim MVN(0,(\tau Q)^{-1})$, $\tau$ is the precision parameter with Gamma prior $Gamma(0.001,0.001)$ and $Q$ is the corresponding inverse covariance kernel of a random walk with boundary correction. For more technical details, see Palacios and Minin (2012). The integral in the exponent of Eq. 3 is then approximated by its Riemann sum. We adapted the implementation for binary trees in phylodyn (Karcher et al., 2017) to the MMC with likelihood Eq. 3. Given $\boldsymbol{m,s,t,n}$, and $\alpha$, our implemented R function BNPR_Lambda() estimates posterior mean and 95% BCI of $N_{e}(t)$ by an Integrated Nested Laplace Approximation (INLA) (Rue et al., 2009).